OpenStudy (anonymous):

It takes Gabriella 4 hours to knit 6 scarves. Assuming that the number of scarves made varies directly as the time spent knitting, how many scarves will she make in 6 hours? Please show your direct variation :)

OpenStudy (anonymous):

Here's a start:\[\text{rate}=\frac{6 \text{ scarves}}{4\text{ hours}}=\cdots\]

OpenStudy (anonymous):

9 scarves

OpenStudy (anonymous):

y = kx 6 = 4x x = 3/2 y = 6x) y = 6(3/2) y = 9